OpenStudy (anonymous):

What is the missing exponent? [ ] (2^-5) =2^-15 The missing exponent is in the brackets, I just need help finding it.

OpenStudy (anonymous):

After this one, I have one more question if your willing to help out.

OpenStudy (jhannybean):

Say the missing exponent = x. Then... \[(2^{-5}) (x) = 2^{-15}\] we can find x by dividing the two given exponents. \(\dfrac{2^{-15}}{2^{-5}}\) When dividing exponents, the powers subtract eachother. \(\dfrac{x^m}{x^{n}} = x^{m-n}\)

OpenStudy (jhannybean):

Do you see what I mean?

OpenStudy (anonymous):

Sort of

OpenStudy (anonymous):

Yea ok, I get it now

OpenStudy (jhannybean):

What did yu get as the missing exponent?

OpenStudy (jhannybean):

Okay. Lets solve for \(\dfrac{2^{-15}}{2^{-5}}\) first. What is \(2^{-15+5} =~?\)

OpenStudy (anonymous):

2^-10 also, how do you get the power of numbers without the ^

OpenStudy (jhannybean):

-15 + 5 = -20??

OpenStudy (jhannybean):

Yes, you are right, it is -5, but remember our rule? \(\large \dfrac{2^{-15}}{2^{-5}} = 2^{-15 -(-5)}= 2^{-15 + 5}\)

OpenStudy (anonymous):

oh yea, so it is 2^-10 ok

OpenStudy (jhannybean):

Remember, 2 negatives make a positive :D

OpenStudy (jhannybean):

Mmhmm, it's 2\(^{-10}\)

OpenStudy (jhannybean):

So do you agree that if \(x=2^{-10}\), that \((2^{-5})(2^{-10}) = 2^{-15}\)?

OpenStudy (anonymous):

Yes, so then in the blank, I would put 2^-10

OpenStudy (jhannybean):

Yes :D \(\checkmark\)

OpenStudy (anonymous):

thanks